OpenStudy (anonymous):

What are the solutions of the rational equation? Use a graphing calculator to solve. (5.66/y)+(5.5/4y)=4

OpenStudy (aravindg):

Used a graphing calculator?

OpenStudy (anonymous):

I don't understand these questions either...

OpenStudy (anonymous):

@yummydum

OpenStudy (anonymous):

I cant afford one T_T

OpenStudy (anonymous):

@AravindG

OpenStudy (aravindg):

you could use it online for free

OpenStudy (anonymous):

Where?

OpenStudy (anonymous):

\[{5.66\over y}+{5.5\over4y}=4\]\[{22.64\over4y}+{5.5\over4y}=4\]\[{28.14\over4y}=4\]\[{28.14\over4}=4y\]\[y=...?\]

OpenStudy (aravindg):

Here is a good one : https://www.desmos.com/calculator

OpenStudy (anonymous):

oh wait..i dont even know if that's right..

OpenStudy (anonymous):

How so?

OpenStudy (anonymous):

err...@AravindG i graphed it on a calculator and im quite confused...it doesnt touch the x-axis :S ..

OpenStudy (anonymous):

All alone...

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

That online calculator wont work '_' .

OpenStudy (whpalmer4):

\[\frac{5.66}{y}+\frac{5.5}{4y}=4\]Multiply through by \(4y\) \[4*5.66+5.5=16y\]Divide by 16\[28.14/16 = y\]\[y=1.75875\]

OpenStudy (anonymous):

So you would replace the y var with 1.75?

OpenStudy (anonymous):

i did exactly what whpalmer4 did, and that is correct, shouldnt have doubted myself lol, the calculator part though, does not work

OpenStudy (anonymous):

@yummydum @whpalmer4, I can always count on you guys at the end.

OpenStudy (whpalmer4):

The solution is the line y = 1.75875...which intersects one of the hyperbolas in the 1st quadrant

OpenStudy (anonymous):

glad i could help

OpenStudy (anonymous):

Could it possibly be negative though?

OpenStudy (anonymous):

-1.75 is one of the choices

OpenStudy (anonymous):

as far as i see, the answer should be positive

OpenStudy (anonymous):

Thank you c: .